# The EMPI Code for Plasma-Induced Effects on Radio Waves I: Non-Magnetized Media and Applications to Fast Radio Bursts

### v1

@@ Abstract
Electromagnetic waves undergo modifications as they propagate through plasma. We present EMPI (ElectroMagnetic-wave Plasma Interaction), a three-dimensional numerical framework designed to simulate the interaction between radio signals and cold plasma. With input plasma density profiles, intrinsic radio signals, and the time and frequency resolutions of the telescope, the code synthesizes observed signals using first-principles calculations. EMPI is capable of modeling a wide range of plasma distributions, spanning analytically described smooth functions (e.g., Gaussian or exponential profiles), statistical models (e.g., turbulent screens), and discrete macroscopic structures like isolated plasma clumps, which are difficult to model both analytically and statistically. Validation tests demonstrate excellent agreement with established plasma propagation effects, such as dispersion, lensing, scintillation, and scattering. This code provides an efficient method for handling both analytical and statistical scenarios, bridging the gap between these descriptions. Thanks to its comprehensive capabilities, EMPI is particularly useful for studying radio sources with cosmological origin, especially pulse-like signals such as Fast Radio Bursts (FRBs). As these signals travel through diverse and complex plasma environments across the universe, their properties are inevitably altered, resulting in observable changes. In this context, EMPI serves as a valuable tool for studying the propagation effects of these sources, helping to advance the understanding of their essence and the intervening plasma environments.
@@ 1 Introduction
Fast radio bursts (FRBs) emerged as a key focus in contemporary astronomy since the discovery of the first FRB exhibiting excess dispersion measure (DM) (Lorimer et al., 2007). Subsequent studies (Thornton et al., 2013; Tendulkar et al., 2017) have confirmed their extragalactic origins, establishing the sources at cosmological distances. Propagating over such enormous distances, these signals inevitably traverse diverse plasma media that modify their observable characteristics. Understanding these propagation effects that cause deviations from the intrinsic characteristics of the bursts is crucial both for advancing the knowledge of FRB mechanisms which remain under debate (Zhang, 2023) and for probing the nature of intervening plasma (Zheng et al., 2014).
@@
Numerous studies have systematically explored the impact of plasma on FRB signals, successfully accounting for key propagation phenomena including dispersion (McQuinn, 2013; Yang & Zhang, 2017; Kulkarni, 2020; Macquart et al., 2020), Faraday rotation and conversion (Feng et al., 2022; Qu & Zhang, 2023; Yang et al., 2023; Xia et al., 2023), plasma lensing (Clegg et al., 1998; Cordes et al., 2017; Er et al., 2020), scintillation (Rickett, 1977, 1990; Kumar et al., 2023) and scattering (Luan & Goldreich, 2014; Cordes et al., 2016; Xu & Zhang, 2016). Building on these foundations, recent observational and theoretical advances have further expanded our understanding of these effects. Observations unveiled novel signatures: wideband propagation effects (Gopinath et al., 2023), precise timescale measurements (Main et al., 2021), scattering variability (Ocker et al., 2022), and scintillation arc morphologies (Wu et al., 2024, 2023). Concurrently, refined propagation models now quantitatively explain FRB-related phenomena, from burst enhancement or suppression (Chen et al., 2024), scattering-induced pulse broadening (Wei et al., 2024), to spectral changes (Levkov et al., 2022; Kumar et al., 2024). Such theoretical progress has transformed propagation effects into diagnostic tools for both sight line plasma characterization (Prochaska & Zheng, 2019; Shin et al., 2024) and radiation mechanism constraints (Kumar et al., 2023; Nimmo et al., 2024). This evolution from qualitative estimation to quantitative modeling (Sammons et al., 2023; Main et al., 2023) enables more precise interpretations. However, it exposed critical limitations: Current approaches predominantly rely on idealized plasma configurations, i.e., whether analytic density profiles (e.g., Gaussian/exponential lenses) or statistical turbulence (e.g., Kolmogorov-type screens). Such simplifications starkly contrast with realistic plasma structures exhibiting unmodeled complexity in both spatial configuration and temporal evolution. Where analytical descriptions fail and statistical assumptions break down (configurations lacking analytical descriptions and sufficient randomness for statistical treatment), elaborate numerical modelings become imperative for accurate quantification.
@@
To bridge this gap, we present EMPI (ElectroMagnetic-wave Plasma Interaction), a three-dimensional (3D) numerical framework based on first principles enabling calculations of electromagnetic wave propagation through arbitrarily structured plasma media. Validated against established theories, EMPI holistically captures comprehensive propagation effects rather than isolating individual phenomena, enabling precise modeling of non-canonical plasma geometries. This versatility facilitates critical constraints on plasma properties imprinting on FRB signals during cosmic traversal.
@@
Crucially, the urgency for such advanced modeling tools is amplified by impending observational revolutions. The Five-hundred-meter Aperture Spherical Radio Telescope (FAST), Square Kilometre Array (SKA), and other large-scale facilities are poised to detect fainter bursts and subtle plasma imprints and map dynamic structural evolution. These observational capacities would remain underexploited without corresponding advances in plasma interaction modeling. EMPI serves as an essential tool for investigating the propagation effects of FRBs, supporting future studies on novel observations. Beyond FRBs, the framework extends naturally to plasma-induced effects studies of other radio emitters such as pulsars and persistent radio sources.
@@
As the inaugural work in the EMPI code series, this paper establishes the foundational framework for non-magnetized plasma regimes, comprehensively addressing all key propagation effects: dispersion, plasma lensing, scintillation, and scattering. Deliberately excluding magnetized plasma and relativistic, nonlinear effects (see dedicated discussions in Luan & Goldreich (2014); Kumar & Lu (2020); Yang & Zhang (2020)), this work focuses on developing first-principles solutions for linear propagation effects. Subsequent studies will seamlessly integrate magnetic field effects and explore advanced scientific applications.
@@
This paper is structured as follows: Section 2 provides the theoretical background of plasma interactions with FRBs; Section 3 details the numerical architecture with modular implementation strategies; Section 4 validates the code through (i) benchmark tests against analytical solutions and (ii) demonstrations of complex scenario handling; Section 5 synthesizes the performance of the code while charting future research directions, focusing on (i) refining plasma lensing and scattering models and (ii) developing parameter constraint methods for observational studies.
@@ 2 Code Principles and Methodology
The plasma-FRB interaction is inherently multifaceted. For clarity, this study restricts its scope to non-magnetized plasma regimes, where time delays, phase shifts, and the electron encountered collectively manifest as four dominant effects: dispersion, plasma lensing, scintillation, and scattering. This focus is justified, since non-magnetized plasma interactions not only pervade astrophysical environments but also constitute the necessary foundation for probing intricate magnetized plasma effects.
@@ 2.1 Criteria for Regime Selection
Plasma-induced electromagnetic wave interactions alter the propagation direction of light via two distinct mechanisms: refraction and diffraction. Refraction occurs when plasma structures are significantly larger than the signal wavelength, permitting a geometric optics approximation. In contrast, diffraction dominates for sub-wavelength plasma inhomogeneities, necessitating a wave-optics treatment based on the Fresnel-Kirchhoff integral. A detailed discussion of these regimes is provided in Jow et al. (2023). Fundamentally, the choice between refractive and diffractive treatment hinges on the relative size of the plasma structures compared to the signal wavelength. In this paper, we derive the criterion by comparing the refractive and diffractive deflection angles caused by the same plasma clump, thereby determining which effect is dominant (Draine, 2011) (see Appendix A for further details). The critical size $l_{\rm c}$ is defined as:
@@
where $m_{\rm e}$ is the electron mass, $\nu$ is the frequency, $c$ is the speed of light in the vacuum, $q$ is the electron charge, and $\Delta n$ represents the characteristic electron density fluctuation measured at the scale of the plasma clump. This critical size delineates the transition between refractive and diffractive regimes: when the physical plasma bump size $l$ is much greater than $l_{\rm c}$, refraction dominates, conversely, when $l\ll l_{\rm c}$, diffraction prevails. This criterion provides a fundamental basis for distinguishing the dominant physical mechanism and guides the selection of appropriate computational methodologies tailored to each regime.
@@
The characteristic physical bump size is an intrinsic property of the input plasma profile, theoretically spanning arbitrary magnitudes. However, numerical calculations introduce resolution-dependent constraints. The plasma screen is described by discrete grid points in 3D space, with the density precisely defined at each point. Grid spacing establishes the minimum resolvable feature size. A bump must occupy multiple adjacent nodes to form discernible structures. Consequently, the physical size of a plasma bump cannot be smaller than the resolution.
@@ 2.2 Refractive Regime
Conventionally, geometric optics focuses on light propagation and redistribution by treating light as rays, often neglecting phase information, since it is not crucial for incoherent light. Nevertheless, when applied to radio waves, especially coherent signals, phase information becomes significant. In this method, phase shifts are calculated alongside ray trajectories, allowing geometric optics to not only trace ray paths but also capture phase variations along each trajectory. This integration ensures accurate phase accounting, which is critical for analyzing coherent signals.
Within this framework, the refractive regime represents light as rays emitted from a point source, each following an independent trajectory perpendicular to the local wavefront surface. When light penetrates a plasma screen, spatial density variations induce position-dependent phase shifts, causing distortions in the wavefront surface and the deflection of light. To quantify this phenomenon, the plasma screen is divided into 3D discrete grid points. As light propagates through the screen, phase shifts are determined by the plasma density. The phase shift $\phi$ consists of two primary contributions: the geometric component (first term in Eq. 2), originating from the distance traveled along the ray trajectory, and the dispersive component (second term in Eq. 2), resulting from the interaction with the non-magnetized plasma. This relationship is described by the following equation:
@@
which is derived from the plasma dispersion relation (see Appendix A for details), where $\omega$ is the angular frequency and $\omega_{\rm p}$ is the plasma frequency. The plasma frequency $\omega_{\rm p}$ can be expressed as $\omega_{\rm p}=\sqrt{4\pi q^{2}n_{\rm e}/m_{\rm e}}$, where $n_{\rm e}$ is the number density of electrons. With the phase shift known, the new local wavefront can be obtained, and Principal Component Analysis (PCA) [^1] Principal Component Analysis (PCA) is a statistical method that identifies the directions (or principal components) along which data varies the most. The process involves computing the eigenvectors and eigenvalues of the covariance matrix of the data points. These eigenvectors represent the axes of maximum variance, and the eigenvalue magnitudes indicate the strength of variance along each axis. The axis with the smallest eigenvalue, indicating the least variation, is taken as the normal direction of the wavefront., is used to find the normal vectors, indicating the new propagation direction of the light.
@@
The iteration process is delineated in detail in Figure 2. It begins at the incident point on the plasma screen, where the initial direction of each ray is determined by the vector connecting the source to its intersection point on the screen. Due to the astronomical distance between the source and the screen, these vectors are nearly parallel across different rays, with only small deviations. The incident position and initial direction set the first step of the iteration and provide the starting conditions for subsequent calculations. In the $i$-th step (layer), the incoming direction and position of the light ray are used to calculate the intersection point with the next layer, defining the position of the $(i+1)$-th step (see panel (b) in Figure 2). The distance between these two points $d_{i}$ represents the distance traveled by the light in the $i$-th layer. This distance is subsequently used to calculate travel time $t_{i}$, phase shift $\phi_{i}$, and the number of electrons encountered or dispersion measure ${\rm DM}_{i}$ for this layer.
@@
To compute the new local wavefront at each intersection point, the position and density information from its surrounding $n$ grid points, where the density is precisely defined, are utilized. A minimum of three points is required for the calculation; however, four points are used in this implementation to avoid numerical instabilities. The incident direction allows for the reconstruction of an initial local wavefront plane (orange plane in panel (c) of Figure 2). The 4 surrounding points are projected onto this wavefront plane (orange points in panel (c) of Figure 2). The key is to identify the 4 corresponding new wavefront points. These 4 surrounding points provide 4 distinct phase shift values, based on the same distance $d_{i}$ but different densities, according to Eq. 2. Then the 4 phase shift values are converted into spatial distances using $\Delta_{\phi}=\phi\lambda/2\pi$, where $\lambda$ is the wavelength. The positions of the new wavefront points in 3D space are obtained by taking the old wavefront points on the wavefront plane and adding the 4 corresponding spatial distances $\Delta_{\phi}$ along the $i$-th direction. (see panel (d) in Figure 2). The direction of the normal vector is determined by applying the PCA method to the 4 new wavefront points, which establishes the direction for the $(i+1)$-th step (see panel (e) in Figure 2).
In each step, relevant physical parameters, such as travel time $t_{i}$, phase shift $\phi_{i}$ and dispersion measure ${\rm DM}_{i}$, are computed as well. The density at the intersection point in the $i$-th step $n_{{\rm e},i}$ is obtained by obtained by interpolating the density of the surrounding grid points. Specifically, the travel time $t_{i}$ is given by the following relation:
@@
Upon completion of one iteration cycle, the accumulated time, phase, and dispersion measure (DM) values are stored for subsequent computational stages.
The iterative procedure numerically determines the final propagation vectors and spatial coordinates of the light rays at the plasma screen exit plane. To evaluate observational detectability, the trajectory of each ray is analyzed using the derived exit parameters and the predetermined separation distance $d_{\rm po}$ between the plasma screen and the observer plane. These parameters enable efficient determination of the intersection point of each ray with the observer plane. By comparing these intersection coordinates with the predefined spatial position of the observer, the rays that can be observed are selected directly to construct the received signal in the next step. Rays that do not reach the observer still provide meaningful insights. The observer plane, which is partitioned into discrete bins, contains the intersection point of each ray. The density of intersection points within each bin is computed, yielding a spatial distribution of light intensity. This distribution corresponds to the concept of gain in lensing literature.

### v2

@@ Abstract
Electromagnetic waves undergo modifications as they propagate through plasma. We present electromagnetic wave plasma interaction (EMPI), a three-dimensional numerical framework designed to simulate the interaction between radio signals and cold plasma. With input plasma density profiles, intrinsic radio signals, and the time and frequency resolutions of the telescope, the code synthesizes observed signals using first-principles calculations. EMPI is capable of modeling a wide range of plasma distributions, spanning analytically described smooth functions (e.g., Gaussian or exponential profiles), statistical models (e.g., turbulent screens), and discrete macroscopic structures like isolated plasma clumps, which are difficult to model both analytically and statistically. Validation tests demonstrate excellent agreement with established plasma propagation effects, such as dispersion, lensing, scintillation, and scattering. This code provides an efficient method for handling both analytical and statistical scenarios, bridging the gap between these descriptions. Thanks to its comprehensive capabilities, EMPI is particularly useful for studying radio sources of cosmological origin, especially pulse-like signals such as fast radio bursts. As these signals travel through diverse and complex plasma environments across the universe, their properties are inevitably altered, resulting in observable changes. In this context, EMPI serves as a valuable tool for studying the propagation effects of these sources, helping to advance the understanding of their essence and the intervening plasma environments.
@@ 1 Introduction
Fast radio bursts (FRBs) emerged as a key focus in contemporary astronomy since the discovery of the first FRB exhibiting excess dispersion measure (DM) (Lorimer et al., 2007). Subsequent studies (Thornton et al., 2013; Tendulkar et al., 2017) have confirmed their extragalactic origins, establishing the sources at cosmological distances. Propagating over such enormous distances, these signals inevitably traverse diverse plasma media that modify their observable characteristics. Understanding these propagation effects that cause deviations from the intrinsic characteristics of the bursts is crucial both for advancing the knowledge of FRB mechanisms, which remain under debate (Zhang, 2023), and for probing the nature of intervening plasma (Zheng et al., 2014).
@@
Numerous studies have systematically explored the impact of plasma on FRB signals, successfully accounting for key propagation phenomena including dispersion (McQuinn, 2013; Yang & Zhang, 2017; Kulkarni, 2020; Macquart et al., 2020), Faraday rotation and conversion (Feng et al., 2022; Qu & Zhang, 2023; Yang et al., 2023; Xia et al., 2023; Wang et al., 2025), plasma lensing (Clegg et al., 1998; Cordes et al., 2017; Er et al., 2020), scintillation (Rickett, 1977, 1990; Kumar et al., 2023) and scattering (Luan & Goldreich, 2014; Cordes et al., 2016; Xu & Zhang, 2016). Building on these foundations, recent observational and theoretical advances have further expanded our understanding of these effects. Observations unveiled novel signatures: wideband propagation effects (Gopinath et al., 2023), precise timescale measurements (Main et al., 2021), scattering variability (Ocker et al., 2022), and scintillation arc morphologies (Wu et al., 2024, 2023). Concurrently, refined propagation models now quantitatively explain FRB-related phenomena, from burst enhancement or suppression (Chen et al., 2024), scattering-induced pulse broadening (Wei et al., 2024), to spectral changes (Levkov et al., 2022; Kumar et al., 2024). Such theoretical progress has transformed propagation effects into diagnostic tools for both sight line plasma characterization (Prochaska & Zheng, 2019; Shin et al., 2024) and radiation mechanism constraints (Kumar et al., 2023; Nimmo et al., 2024). This evolution from qualitative estimation to quantitative modeling (Sammons et al., 2023; Main et al., 2023) enables more precise interpretations. However, it exposed critical limitations: Current approaches predominantly rely on idealized plasma configurations, i.e., whether analytic density profiles (e.g., Gaussian/exponential lenses) or statistical turbulence (e.g., Kolmogorov-type screens). Such simplifications starkly contrast with realistic plasma structures exhibiting unmodeled complexity in both spatial configuration and temporal evolution. Where analytical descriptions fail and statistical assumptions break down (configurations lacking analytical descriptions and sufficient randomness for statistical treatment), elaborate numerical modeling becomes imperative for accurate quantification.
@@
To bridge this gap, we present EMPI, a three-dimensional (3D) numerical framework based on first principles, enabling calculations of electromagnetic wave propagation through arbitrarily structured plasma media. Validated against established theories, EMPI holistically captures comprehensive propagation effects rather than isolating individual phenomena, enabling precise modeling of noncanonical plasma geometries. This versatility facilitates critical constraints on plasma properties imprinting on FRB signals during cosmic traversal.
@@
Crucially, the urgency for such advanced modeling tools is amplified by impending observational revolutions. The Five hundred-meter Aperture Spherical Radio Telescope, Square Kilometre Array, and other large-scale facilities are poised to detect fainter bursts and subtle plasma imprints and map dynamic structural evolution. These observational capacities would remain underexploited without corresponding advances in plasma interaction modeling. EMPI serves as an essential tool for investigating the propagation effects of FRBs, supporting future studies on novel observations. Beyond FRBs, the framework extends naturally to plasma-induced effects studies of other radio emitters such as pulsars and persistent radio sources.
@@
As the inaugural work in the EMPI code series, this paper establishes the foundational framework for nonmagnetized plasma regimes, comprehensively addressing all key propagation effects: dispersion, plasma lensing, scintillation, and scattering. Deliberately excluding magnetized plasma and relativistic, nonlinear effects (see dedicated discussions in Luan & Goldreich (2014); Kumar & Lu (2020); Yang & Zhang (2020)), this work focuses on developing first-principles solutions for linear propagation effects. Subsequent studies will seamlessly integrate magnetic field effects and explore advanced scientific applications.
@@
This paper is structured as follows: Section 2 provides the theoretical background of plasma interactions with FRBs; Section 3 details the numerical architecture with modular implementation strategies; Section 4 validates the code through (i) benchmark tests against analytical solutions and (ii) demonstrations of complex scenario handling; Section 5 summarizes the performance of the code while charting future research directions, focusing on (i) refining plasma lensing and scattering models and (ii) developing parameter constraint methods for observational studies.
@@ 2 Code Principles and Methodology
The plasma–FRB interaction is inherently multifaceted. For clarity, this study restricts its scope to nonmagnetized plasma regimes, where time delays, phase shifts, and the electron encountered collectively manifest as four dominant effects: dispersion, plasma lensing, scintillation, and scattering. This focus is justified, since nonmagnetized plasma interactions not only pervade astrophysical environments but also constitute the necessary foundation for probing intricate magnetized plasma effects.
@@ 2.1 Criteria for Regime Selection
Plasma-induced plasma-EM wave interactions alter the propagation direction of light via two distinct mechanisms: refraction and diffraction. Refraction occurs when plasma structures are significantly larger than the signal wavelength, permitting a geometric optics approximation. In contrast, diffraction dominates for sub-wavelength plasma inhomogeneities, necessitating a wave-optics treatment based on the Fresnel–Kirchhoff integral. A detailed discussion of these regimes is provided in Jow et al. (2023). Fundamentally, the choice between refractive and diffractive treatment hinges on the relative size of the plasma structures compared to the signal wavelength. In this paper, we derive the criterion by comparing the refractive and diffractive deflection angles caused by the same plasma clump, thereby determining which effect is dominant (Draine, 2011) (see Appendix A for further details). The critical size $l_{\rm c}$ is defined as:
@@
where $m_{\rm e}$ is the electron mass, $\nu$ is the frequency, $c$ is the speed of light in the vacuum, $q$ is the electron charge, and $\Delta n$ represents the characteristic electron density fluctuation measured at the scale of the plasma clump. This critical size delineates the transition between refractive and diffractive regimes: when the physical plasma bump size $l$ is much greater than $l_{\rm c}$, refraction dominates, conversely, when $l\ll l_{\rm c}$, diffraction prevails. This criterion establishes theoretical foundations for distinguishing dominant physical mechanism and informs the selection of appropriate computational methodologies tailored to distinct regimes.
@@
The characteristic physical bump size is an intrinsic property of the input plasma profile, theoretically spanning arbitrary magnitudes. However, numerical calculations introduce resolution-dependent constraints. The plasma screen is described by discrete grid points in 3D space, with the density precisely defined at each point. Grid spacing establishes the minimum resolvable feature size. A bump must occupy multiple adjacent nodes to form discernible structures. Consequently, the physical size of a plasma bump cannot be smaller than the grid resolution.
@@ 2.2 Refractive Regime
In conventional geometric optics, the ray-based approach to studying light propagation and redistribution inherently neglects phase information, as it is not crucial for incoherent light. Nevertheless, when applied to radio waves, especially coherent signals, phase information becomes significant. In this method, phase shifts are calculated alongside ray trajectories, allowing geometric optics to not only trace ray paths but also capture phase variations along each trajectory. This integration ensures accurate phase accounting, which is essential for analyzing coherent signals.
Within this framework, the refractive regime models light as rays originating from a point source, where each ray propagates independently along a trajectory perpendicular to the local wave front surface. When light propagates through a plasma screen, spatial variations in the plasma density generate position-dependent phase shifts. These shifts distort the wave front surface and induce deflections in the light propagation direction. To computationally quantify this effect, the plasma screen is divided into 3D discrete grid points. As the light penetrates the screen, phase shifts are determined by the plasma density.
The phase shift $\phi$ consists of two primary contributions: the geometric component (first term in Eq. 2), originating from the distance traveled along the ray trajectory, and the dispersive component (second term in Eq. 2), resulting from the interaction with the non-magnetized plasma. This relationship is described by the following equation:
@@
which is derived from the plasma dispersion relation (see Appendix A for details), where $\omega$ is the angular frequency and $\omega_{\rm p}$ is the plasma frequency. The plasma frequency $\omega_{\rm p}$ can be expressed as $\omega_{\rm p}=\sqrt{4\pi q^{2}n_{\rm e}/m_{\rm e}}$, where $n_{\rm e}$ is the number density of electrons. With the phase shift known, the new local wave front can be obtained, and principal component analysis (PCA) [^1] PCA is a statistical method that identifies the directions (or principal components) along which data varies the most. The process involves computing the eigenvectors and eigenvalues of the covariance matrix of the data points. These eigenvectors represent the axes of maximum variance, and the eigenvalue magnitudes indicate the strength of variance along each axis. The axis with the smallest eigenvalue, indicating the least variation, is taken as the normal direction of the wave front. is used to find the normal vectors, indicating the new propagation direction of the light.
@@
The iteration process is delineated in detail in Figure 2. It begins at the incident point on the plasma screen, where the initial direction of each ray is determined by the vector connecting the source to its intersection point on the screen. Due to the astronomical distance between the source and the screen, these vectors are nearly parallel across different rays, with only small deviations. The incident position and initial direction set the first step of the iteration and provide the starting conditions for subsequent calculations. In the $i$-th step (layer), the incoming direction and position of the light ray are used to calculate the intersection point with the next layer, defining the position of the ($i$+1)-th step (see panel (b) in Figure 2). The distance between these two points $d_{i}$ represents the distance traveled by the light in the $i$-th layer. This distance is subsequently used to calculate travel time $t_{i}$, phase shift $\phi_{i}$, and the number of electrons encountered or dispersion measure ${\rm DM}_{i}$, for this layer.
@@
To compute the updated local wave front at each intersection point, the position and density information from its $n$ surrounding grid points, where the density is precisely defined, are utilized. A minimum of three points is required for the calculation; however, four points are used in this implementation to avoid numerical instabilities. The incident direction allows for the reconstruction of an initial local wave front plane (orange plane in panel (c) of Figure 2). The four surrounding points are projected onto this wave front plane (orange points in panel (c) of Figure 2). The key is to identify the four corresponding new wavefront points. These four surrounding points yield four distinct phase shift values, calculated from identical distances $d_{i}$ but different densities, via Eq. 2. Then the four phase shift values are converted into spatial displacements using $\Delta_{\phi}=\phi\lambda/2\pi$, where $\lambda$ is the wavelength. The positions of the updated wave front points in 3D space are obtained by offsetting the initial wave front points on the original wave front plane by their respective spatial distances $\Delta_{\phi}$ along the $i$-th direction. (see panel (d) in Figure 2). Finally, the normal vector direction of the updated local wave front is determined by applying the PCA method to these four updated wave front points, defining the propagation direction for the ($i$+1)-th step (see panel (e) in Figure 2).
In each step, relevant physical parameters, such as travel time $t_{i}$, phase shift $\phi_{i}$ and dispersion measure ${\rm DM}_{i}$, are computed as well. The density at the intersection point in the $i$-th step $n_{{\rm e},i}$ is obtained by interpolating the density of the surrounding grid points. Specifically, the travel time $t_{i}$ is given by the following relation:
@@
Upon completion of one iteration cycle, the accumulated time, phase, and DM values are stored for subsequent computational stages.
The iterative procedure numerically determines the final propagation vectors and spatial coordinates of the light rays at the plasma screen exit plane. To evaluate observational detectability, the trajectory of each ray is analyzed using the derived exit parameters and the predefined separation distance $d_{\rm po}$ between the plasma screen and the observer plane. These parameters enable direct determination of the intersection point of each ray with the observer plane. By comparing these intersection coordinates with the predefined spatial position of the observer, the rays that can be observed are selected to construct the received signal in the next step. Rays that do not reach the observer still provide meaningful insights. The observer plane, which is partitioned into discrete bins, contains the intersection point of each ray. The density of intersection points within each bin is computed, yielding a spatial distribution of light intensity. This distribution corresponds to the concept of gain in lensing literature.